OpenStudy (anonymous):

How many days will it take for a sum of $1100 to earn $24 interest if it is deposited in a bank paying 6%/year? (Use a 365-day year. Round your answer up to the nearest full day.)

OpenStudy (anonymous):

How is it compounded?

OpenStudy (anonymous):

doesnt say

OpenStudy (anonymous):

I will assume it is daily

OpenStudy (anonymous):

yea that is what I am thinking

OpenStudy (anonymous):

1100-Principle(1+Rate/N)^(N)=24

OpenStudy (anonymous):

Try plugging in

OpenStudy (anonymous):

ok thanks

OpenStudy (gw2011):

I got an answer of 133 days. This is how I calculated it: I=(P)(R)(T) I=(1100)(0.06)(1) I=$66 Since the interest asked for is $24, then this will be a fraction of one year. Therefore, the following calculation is done: (24/66)(365)=132.7 days=133 days rounded to the nearest day

OpenStudy (anonymous):

that is correct!!! thank you!!

OpenStudy (gw2011):

You're welcome.